# Smooth Lower Bounds for Differentially Private Algorithms via Padding-and-Permuting Fingerprinting Codes

### v3

@@ Abstract
Fingerprinting arguments, first introduced by [BUV14, ], are the most widely used method for establishing lower bounds on the sample complexity or error of approximately differentially private (DP) algorithms. Still, there are many problems in differential privacy for which we don’t know suitable lower bounds, and even for problems that we do, the lower bounds are not smooth, and become vacuous when the error is larger than some threshold.
@@
    A tight lower bound for DP averaging in the low-accuracy regime, which in particular implies a lower bound for the private $1$-cluster problem introduced by [NSV16, ].
@@
    A lower bound on the additive error of DP algorithms for approximate $k$-means clustering, as a function of the multiplicative error, which is tight for a constant multiplication error.
@@
    A lower bound for estimating the top singular vector of a matrix under DP in low-accuracy regimes, which is a special case of the DP subspace estimation problem studied by [SS21a, ].
Our new tools are based on applying a padding-and-permuting transformation to a fingerprinting code. However, rather than proving our results using a black-box access to an existing fingerprinting code (e.g., Tardos’ code [Tar08]), we develop a new fingerprinting lemma that is stronger than those of [DSS+15, BSU17, ], and prove our lower bounds directly from the lemma. Our lemma, in particular, gives a simpler fingerprinting code construction with optimal rate (up to polylogarithmic factors) that is of independent interest.
@@ 1 Introduction
Differentially private (DP) [DMNS06] algorithms provide a strong guarantee of privacy to the individuals who contribute their data. Informally, a DP algorithm takes data from many individuals and guarantees that no attacker, regardless of their knowledge or capabilities, can learn much more about any one individual than they would have if that individual’s data had never been collected. There is a large body of literature on DP algorithms, and DP algorithms have now been deployed by large technology companies and government organizations.
@@
However, in many applications of DP involving real data, we do not want to assume that the center of the ball is known. For example, algorithms like clustering, covariance estimation, and PCA are often applied to datasets as a preprocessing step to understand the general properties, and we cannot assume the user already knows the location of the data. Thus, we want to assume that the data lies in a ball whose center is unknown. For this problem, the FriendlyCore algorithm of [TCK+22, ] is able to achieve the same error guarantee even when the location of the data is a priori unknown, provided $n\gtrsim\sqrt{d}/\lambda$.
Lower Bounds for DP Averaging.    The work of [BUV14, ] proved that $\Omega(\sqrt{d}/\lambda)$ samples are required for DP averaging, when $\lambda\lesssim 1$. Their work introduced the method of fingerprinting codes to differential privacy, and this technique has become the standard approach for proving lower bounds for differentially private algorithms, either by reduction to the averaging problem or by non-black-box use of the fingerprinting technique (see Related Work).
The lower bound of [BUV14, ] has a significant drawback that it only applies when $\lambda\lesssim 1$, and the lower bound on the sample complexity is vacuous for $\lambda\geq 1$. This limitation is inherent to the way these lower bounds work, since they construct a hard distribution over the hypercube $\{-1,1\}^{d}$, which lies in a known ball of radius $r=\sqrt{d}$. So the DP algorithm that outputs $\hat{x}=\vec{0}$ satisfies
@@ 1.1.1 Main Hardness Results
Theorem 1.1 can be proven by combining our padding-and-permuting technique with an optimal fingerprinting code—such as Tardos’ code [Tar08]—in a black-box way. Moreover, in contrast with most recent constructions of DP lower bounds that use only a so-called fingerprinting lemma, our techniques seem to require the using of a fingerprinting code. Specifically, although fingerprinting lemmas are simpler and more flexible, they require a stronger notion of accuracy that does not fit our padding-and-permuting construction, whereas fingerprinting codes require only an extremely weak notion of accuracy to obtain hardness. In an effort to unify and simplify the techniques used to prove DP lower bounds, we give an alternative proof that makes use of a new fingerprinting lemma that only requires very weak accuracy to obtain hardness (see Section 2 for more details).
@@ 1.1.2 Application: Averaging and $1$-Cluster
We remark that for these specific tasks of DP-averaging/1-cluster, a recent work of [NME22], which provides a lower bound for user-level DP averaging, can also be used to prove a similar statement to Theorem 1.3 (up to poly-logarithmic terms). Their technique is very different from ours, and in particular, do not extend to proving Theorems 1.1 and 1.2 which serve as simple tools for proving lower bounds for the other problems. We refer to Section 2.4 for a more detailed comparison.
1.1.3 Application: $k$-Means Clustering
@@
${\rm COST}_{{\cal S}}(C):=\sum_{x\in{\cal S}}\min_{i\in[k]}\mathopen{}\mathclose{{}\left\lVert x-c_{i}}\right\rVert_{2}^{2}.$
@@
${\mathrm{Pr}}_{C\sim\mathsf{M}({\cal S})}\mathopen{}\mathclose{{}\left[{\rm COST}_{{\cal S}}(C)\leq\lambda\cdot{\rm OPT}_{k}({\cal S})+\xi}\right]\geq\beta,$
where ${\rm OPT}_{k}({\cal S}):=\min_{C\in({\cal B}_{d})^{k}}{\rm COST}_{{\cal S}}(C)$.
While non-private algorithms usually do not have an additive error, under DP an additive error is necessary. As far as we are aware, the only known lower bound on the additive error is the one of [GLM+10] (stated for $k$-medians), which has been extended later by [NCX21] (Theorem 1.2). This lower bound essentially says that the additive error $\xi$ of any $(1,\frac{1}{n^{2}})$-DP algorithm for $k$-means (regardless of its multiplicative error) must be at least $\tilde{\Omega}(k)$. However, as far as we are aware, all known DP upper bounds have an additive error of at least $\Omega(k\sqrt{d})$. In particular, this is also the situation in the state-of-the-art upper bounds that have constant multiplicative error.
@@
Furthermore, an additive error of $\tilde{\Theta}(k\sqrt{d})$ also appears in algorithms that provide utility only for datasets that are well-separated into $k$-clusters: For a small parameter $\phi\in[0,1]$, a dataset ${\cal S}\in({\cal B}^{d})^{n}$ is called $\phi$-separated for $k$-means if ${\rm OPT}_{k}({\cal S})\leq\phi^{2}\cdot{\rm OPT}_{k-1}({\cal S})$ ([ORSS12]).
@@
Let $n,k,d\in{\mathbb{N}}$, $\lambda\geq 1$, $\beta\in(0,1]$ and $\xi\geq 0$ such that $n\geq\xi+k$. If $\mathsf{M}\colon({\cal B}_{d})^{n}\rightarrow({\cal B}_{d})^{k}$ is an $\mathopen{}\mathclose{{}\left(1,\frac{\beta}{4nk}}\right)$-DP $(\lambda,\xi,\beta)$-approximation algorithm for $k$-means, then either $k\geq 2^{\Omega(d/\lambda)}\beta\lambda/d$ or $\xi\geq\Omega\mathopen{}\mathclose{{}\left(\frac{k\sqrt{d/\lambda}}{\log^{1.5}\mathopen{}\mathclose{{}\left(\frac{kd}{\beta\lambda}}\right)}}\right)$.
@@ 1.1.4 Application: Top Singular Vector
[DTTZ14, ] proved a general lower bound of $n=\tilde{\Omega}(\sqrt{d})$ for any algorithm that identifies a useful approximation to the top singular vector. Yet, [SS21, ] bypassed this lower bound under the assumption that the points are close to lying in a $1$-dimensional subspace, defined by having a small ratio $\sigma_{2}/\sigma_{1}$. They showed that when the points are Gaussian, and the above ratio is small, the sample complexity can be made independent of $d$. More specifically, their result shows that if $\sigma_{2}/\sigma_{1}\leq\tilde{\Theta}(1/d)$, then $O_{\varepsilon,\delta}(1)$ samples suffice to approximate the top singular vector. This result was later extended by [KMS+22, ] who showed that we can actually achieve a sample complexity that is proportional to $\sigma_{2}/\sigma_{1}$. These works also consider the more general problem of estimating the span of the top $k$ singular vectors, which we do not consider in this work.
@@
For comparison, [DTTZ14, ] proved a lower bound of $\tilde{\Omega}(\sqrt{d})$ on the sample complexity but it only applies when $\gamma$ is larger than some specific constant, whereas our lower bound holds for the entire range of $\gamma$.
@@ 1.2 Related Work
The connection between fingerprinting codes and differential privacy was first introduced by Bun, Ullman, and Vadhan [BUV14]. Subsequent work significantly simplified and generalized the method in a number of ways [SU16, DSS+15, BSU17, SU17, KLSU19, CWZ21, KMS22, CWZ21, NME22], including the removal of fingerprinting codes and distilling the main technical component into a fingerprinting lemma. The price of this simplicity and generality is that the lower bounds rely on having a stronger accuracy requirement that constrains both marked and unmarked columns, whereas the hard instances constructed directly from fingerprinting codes require only a very weak accuracy requirement that constrains only marked columns.
Fingerprinting lower bounds have applied in many settings, including mean estimation [BUV14, DSS+15, KLSU19, NME22], adaptive data analysis [HU14, SU15], empirical risk minimization [BST14], spectral estimation [DTTZ14], combining public and private data [BCM+20], regression [CWZ21, CWZ23], sampling [RSSS21], Gaussian covariance estimation [KMS22, Nar24], continual observation [JRSS23], and unbiased private estimation [KMR+23]. Of these works, all of them except for [RSSS21] use fingerprinting lemmas, rather than the stronger construction of fingerprinting codes. Except [NME22] (discussed in Section 2.4), all these lower bounds are based on the same type of hard instances that lead to the data being contained in a known ball of a given radius, and thus none of them have the smoothness property we desire. Thus, we believe that our method could find other applications beyond the ones that are described in this paper.
@@ 2 Our Technique
In this section, we present a proof overview of Theorem 1.1 for $\beta\approx 1$. In Section 2.1 we present a simple variant of the proof that uses an optimal fingerprinting code as black-box (e.g., [Tar08, ]). In Section 2.2 we explain how we actually avoid the use of Tardos’ fingerprinting code by developing a strong fingerprinting lemma. In Section 2.3 we describe our more general framework, and in Section 2.4 we make a detailed comparison with [NME22].
@@ 2.1 Proof via an Optimal Fingerprinting Code
Fingerprinting codes were originally introduced by [BS98, ]. [Tar08, ] constructed an optimal FPC of length $d_{0}=\tilde{\Theta}(n^{2})$, and [BUV14, ] proved that Tardos’ code is actually robust, i.e., it enables tracing even when the adversary is allowed to be inconsistent with a small fraction of marked columns (say, $20\%$).111[BUV14] did not try to optimize the constant in the fraction of errors, and only proved it for $4\%$. For the purpose of this proof sketch, we assume that the code is robust for $20\%$ errors. A formal proof that relies on their result must change the constant $90\%$ in Section 1.1.1 to a constant larger than $96\%$.
@@
Input parameters: Number of users $n\in{\mathbb{N}}$ and accuracy parameter $\alpha\in[0,1]$. Let $d_{0}=d_{0}(n)$ be the codewords’ length of $\mathsf{Gen}(n)$ and let $d=d_{0}+2\ell$ for $\ell=\mathopen{}\mathclose{{}\left\lceil\frac{d_{0}}{2\alpha}}\right\rceil$.
Operation:
  1. 1.
    Sample a codebook $X\in\{-1,1\}^{n\times d_{0}}$ along with a secret state ${\sf st}$ according to $\mathsf{Gen}(n)$.
  2. 2.
    Append $\ell$ $1$-marked and $\ell$ $(-1)$-marked columns to the matrix $X$.
  3. 3.
    Permute the columns of $X$ according to a random permutation $\pi$ over $[d]$.
  4. 4.
    Output the resulting matrix $X^{\prime}\in\{-1,1\}^{n\times d}$ along with the new state ${\sf st}^{\prime}=(\pi,{\sf st})$.
@@
Input parameters: A weakly-accurate result $q=(q^{1},\ldots,q^{d})\in\{-1,1\}^{d}$ and a secret state ${\sf st}^{\prime}=(\pi,{\sf st})$.
Operation:
  • Output $\mathsf{Trace}(\tilde{q},{\sf st})$ for $\tilde{q}=(q^{\pi(1)},\ldots,q^{\pi(d_{0})})\in\{-1,1\}^{d_{0}}$.
@@
Let $(\mathsf{Gen},\mathsf{Trace})$ be a robust FPC with codewords’ length $d_{0}=\tilde{\Theta}(n^{2})$ (e.g., [Tar08]). Suppose that we sample $X^{\prime}=(x_{1}^{\prime},\ldots,x_{n}^{\prime})\in\{-1,1\}^{n\times d}$ according to $\mathsf{Gen}^{\prime}(n,\alpha)$. By construction, $X^{\prime}$ contains at least $\frac{1}{2}(1-\alpha)d$ $b$-marked columns, for both $b\in\{-1,1\}$. Therefore, if $\mathsf{M}\colon\{-1,1\}^{n\times d}\rightarrow\{-1,1\}^{d}$ is $\alpha$-weakly-accurate, the output $q\in\{-1,1\}^{d}$ of $\mathsf{M}(X^{\prime})$ must agree with $90\%$ of the marked columns of $X^{\prime}$. But because the columns are randomly permuted, $\mathsf{M}$ cannot distinguish between marked columns from the padding and marked columns from the original codebook $X$. This means that it must agree with a similar fraction of marked columns of the codebook $X$, which enables tracing since the code is robust. Therefore, we conclude that such a weakly-accurate mechanism cannot be DP unless $n\geq\tilde{\Omega}(\sqrt{d_{0}})$, i.e., $n\geq\tilde{\Omega}(\sqrt{\alpha d})$.
@@ 2.2 Proof via a Strong Fingerprinting Lemma
The disadvantage of using FPC as black-box for DP lower bounds is the fact that [Tar08]’s analysis is quite involved, so it is hard to gain an end-to-end understanding of the process, and in particular, to generate hard instances in more complicated settings (e.g., exponential families [KMS22]). Therefore, later results simplified the construction and especially the analysis, at the cost of considering more restricted adversaries that must estimate the average of most coordinates, and not just the “marked” ones (which usually suffices for the aggregation tasks we are interested in). This led to the development of the Fingerprinting Lemma (described below) which serves as the most common tool for proving lower bounds for approximate DP algorithms.
@@
Now the same approach of [BSU17, DSS+15] yields that we can increase the correlation by taking $d=\tilde{\Theta}(n^{2})$ independent columns, [^2] An additional factor of $\log^{2}n$ is hidden inside the $\tilde{\Theta}$ due to the $\log n$ factor that we lose in our new fingerprinting lemma. However, we ignore it for the sake of this presentation as it is only a low-order term. which leads to a tracing algorithm. But unlike [BSU17, DSS+15], we obtain a much stronger “FPC-style” tracing algorithm, which enables to apply a similar PAP approach as described in Section 2.1. As a corollary that is of independent interest, the above approach results in a new fingerprinting code that has a simpler analysis than the one of [Tar08], described in Appendix A.
@@
Note that strongly-correlated (Section 2.2) is similar to strongly-agrees (Section 1.1.1) but not exactly the same (in particular, the former is a property of a random variable while the latter is a property of a fixed vector). In our analysis we use the strongly-correlated definition which makes the statements and proofs much more clean. To see this, recall that in the proof sketch in Section 2.1, we claimed that an algorithm that agrees with $90\%$ of the marked columns of $X^{\prime}$, also agrees with a “similar” fraction of marked columns of the original codebook $X$. But this is not exactly the same agreement fraction, and it creates some additional restrictions on the parameters, and in particular, one has to go into the specific FPC construction to argue that w.h.p., there are many marked columns (as done by [BUV14] w.r.t. [Tar08]’s code). But in this work we observed that padding-and-permuting simply transforms a $90\%$ agreement on $X^{\prime}$ (i.e., strong-agreement) into $0.9$-correlation on $X$ (i.e., strong-correlation) without any special requirements on the parameters (this is the content of Section 6), so we were able use our FPC directly.

### v4

@@ Abstract
[…]
@@
[…]
@@
    A lower bound on the additive error of DP algorithms for approximate $k$-means clustering and general $(k,z)$-clustering, as a function of the multiplicative error, which is tight for a constant multiplication error.
@@
[…]
@@ 1 Introduction
[…]
@@
[…]
@@ 1.1.1 Main Hardness Results
[…]
@@ 1.1.2 Application: Averaging and $1$-Cluster
We remark that for these specific tasks of DP-averaging/1-cluster, a recent work of [NarayananME22], which provides a lower bound for user-level DP averaging, can also be used to prove a similar statement to Theorem 1.3 (up to poly-logarithmic terms). Their technique is very different from ours, and in particular, do not extend to proving Theorems 1.1 and 1.2 which serve as simple tools for proving lower bounds for the other problems. We refer to Section 2.4 for a more detailed comparison.
1.1.3 Application: Clustering
@@
${\rm COST}(C;{\cal S}):=\sum_{x\in{\cal S}}\min_{i\in[k]}\mathopen{}\mathclose{{}\left\lVert x-c_{i}}\right\rVert_{2}^{2}.$
@@
${\mathrm{Pr}}_{C\sim\mathsf{M}({\cal S})}\mathopen{}\mathclose{{}\left[{\rm COST}(C;{\cal S})\leq\lambda\cdot{\rm OPT}_{k}({\cal S})+\xi}\right]\geq\beta,$
where ${\rm OPT}_{k}({\cal S}):=\min_{C\in({\cal B}_{d})^{k}}{\rm COST}(C;{\cal S})$.
While non-private algorithms usually do not have an additive error, under DP an additive error is necessary. As far as we are aware, the only known lower bound on the additive error is the one of [GuptaLMRT10] (stated for $k$-medians), which has been extended later by [NguyenCX21] (Theorem 1.2). This lower bound essentially says that the additive error $\xi$ of any $(1,\frac{1}{n^{2}})$-DP algorithm for $k$-means (regardless of its multiplicative error) must be at least $\tilde{\Omega}(k)$. However, as far as we are aware, all known DP upper bounds have an additive error of at least $\Omega(k\sqrt{d})$. In particular, this is also the situation in the state-of-the-art upper bounds that have constant multiplicative error.
@@
Furthermore, an additive error of $\tilde{\Theta}(k\sqrt{d})$ also appears in algorithms that provide utility only for datasets that are well-separated into $k$-clusters: For a small parameter $\phi\in[0,1]$, a dataset ${\cal S}\in({\cal B}^{d})^{n}$ is called $\phi$-separated for $k$-means if ${\rm OPT}_{k}({\cal S})\leq\phi^{2}\cdot{\rm OPT}_{k-1}({\cal S})$ ([OstrovskyRSS12]).
@@
Let $n,k,d\in{\mathbb{N}}$, $\lambda\geq 1$, $\beta\in(0,1]$ and $\xi\geq 0$ such that $n\geq k+80\xi$. If $\mathsf{M}\colon({\cal B}_{d})^{n}\rightarrow({\cal B}_{d})^{k}$ is an $\mathopen{}\mathclose{{}\left(1,\frac{\beta}{4nk}}\right)$-DP $(\lambda,\xi,\beta)$-approximation algorithm for $k$-means, then either $k\geq 2^{\Omega(d/\lambda)}\beta\lambda/d\>$ or $\>\xi\geq\Omega\mathopen{}\mathclose{{}\left(\frac{k\sqrt{d/\lambda}}{\log^{1.5}\mathopen{}\mathclose{{}\left(\frac{kd}{\beta\lambda}}\right)}}\right)$.
@@
We remark that our proof is not tailored to $k$-means clustering. In Section 7.2 we state and prove an extension of Theorem 1.4 to $(k,z)$-clustering in which the cost is measured by the sum of the $z^{\rm th}$ powers of the distances ($k$-means clustering is the special case of $z=2$).
@@ 1.1.4 Application: Top Singular Vector
[…]
DTTZ14 proved a general lower bound of $n=\tilde{\Omega}(\sqrt{d})$ for any algorithm that identifies a useful approximation to the top singular vector. Yet, [SinghalS21] bypassed this lower bound under a distributional assumption which implies that the points are close to lying in a $1$-dimensional subspace, defined by having a small ratio $\sigma_{2}/\sigma_{1}$. They showed that when the points are Gaussian, and the above ratio is small, the sample complexity can be made independent of $d$. They also consider the more general problem of estimating the span of the top $k$ singular vectors, which we do not consider in this work.
@@
[…]
@@ 1.2 Related Work
The connection between fingerprinting codes and differential privacy was first introduced by Bun, Ullman, and Vadhan [BUV14]. Subsequent work significantly simplified and generalized the method in a number of ways [SteinkeU16, DworkSSUV15, BunSU17, SteinkeU17, KLSU19, CaiWZ21, KamathMS22, CaiWZ21, NarayananME22], including the removal of fingerprinting codes and distilling the main technical component into a fingerprinting lemma. The price of this simplicity and generality is that the lower bounds rely on having a stronger accuracy requirement that constrains both marked and unmarked columns, whereas the hard instances constructed directly from fingerprinting codes require only a very weak accuracy requirement that constrains only marked columns.
Fingerprinting lower bounds have applied in many settings, including mean estimation [BUV14, DworkSSUV15, KLSU19, NarayananME22], adaptive data analysis [HU14, SU15], empirical risk minimization [BassilyST14], spectral estimation [DTTZ14], combining public and private data [BassilyCMNUW20], regression [CaiWZ21, CaiWZ23], sampling [RaskhodnikovaSSS21], Gaussian covariance estimation [KamathMS22, Narayanan2024], continual observation [JainRSS23], and unbiased private estimation [KamathMRSSS23]. Of these works, all of them except for [RaskhodnikovaSSS21] use fingerprinting lemmas, rather than the stronger construction of fingerprinting codes. Except [NarayananME22] (discussed in Section 2.4), all these lower bounds are based on the same type of hard instances that lead to the data being contained in a known ball of a given radius, and thus none of them have the smoothness property we desire. Thus, we believe that our method could find other applications beyond the ones that are described in this paper.
@@ 2 Our Technique
[…]
@@ 2.1 Proof via an Optimal Fingerprinting Code
[…]
@@
Input parameters: Number of users $n\in{\mathbb{N}}$ and accuracy parameter $\alpha\in[0,1]$. Let $d_{0}=d_{0}(n)$ be the codewords’ length of $\mathsf{Gen}(n)$ and let $d=d_{0}+2\ell$ for $\ell=\mathopen{}\mathclose{{}\left\lceil\frac{d_{0}}{2\alpha}}\right\rceil$. Operation: 1. Sample a codebook $X\in\{-1,1\}^{n\times d_{0}}$ along with a secret state ${\sf st}$ according to $\mathsf{Gen}(n)$. 2. Append $\ell$ $1$-marked and $\ell$ $(-1)$-marked columns to the matrix $X$. 3. Permute the columns of $X$ according to a random permutation $\pi$ over $[d]$. 4. Output the resulting matrix $X^{\prime}\in\{-1,1\}^{n\times d}$ along with the new state ${\sf st}^{\prime}=(\pi,{\sf st})$.
@@
Input parameters: A weakly-accurate result $q=(q^{1},\ldots,q^{d})\in\{-1,1\}^{d}$ and a secret state ${\sf st}^{\prime}=(\pi,{\sf st})$. Operation: Output $\mathsf{Trace}(\tilde{q},{\sf st})$ for $\tilde{q}=(q^{\pi(1)},\ldots,q^{\pi(d_{0})})\in\{-1,1\}^{d_{0}}$.
Note that we set the padding length as a function of the accuracy parameter $\alpha$ such that a weaker accuracy (i.e., smaller $\alpha$) results with a larger padding. This is crucial for creating hard instances in the regime where an $\alpha$-weakly-accurate mechanism must be accurate.
@@
Let $(\mathsf{Gen},\mathsf{Trace})$ be a robust FPC with codewords’ length $d_{0}=\tilde{\Theta}(n^{2})$ (e.g., [Tardos08]). Suppose that we sample $X^{\prime}=(x_{1}^{\prime},\ldots,x_{n}^{\prime})\in\{-1,1\}^{n\times d}$ according to $\mathsf{Gen}^{\prime}(n,\alpha)$. By construction, $X^{\prime}$ contains at least $\frac{1}{2}(1-\alpha)d$ $b$-marked columns, for both $b\in\{-1,1\}$. Therefore, if $\mathsf{M}\colon\{-1,1\}^{n\times d}\rightarrow\{-1,1\}^{d}$ is $\alpha$-weakly-accurate, then the output $q\in\{-1,1\}^{d}$ of $\mathsf{M}(X^{\prime})$ must agree with $90\%$ of the marked columns of $X^{\prime}$. But because the columns are randomly permuted, $\mathsf{M}$ cannot distinguish between marked columns from the padding and marked columns from the original codebook $X$. This means that it must agree with a similar fraction of marked columns of the codebook $X$, which enables tracing since the code is robust. Therefore, we conclude that such a weakly-accurate mechanism cannot be DP unless $n\geq\tilde{\Omega}(\sqrt{d_{0}})$, i.e., $n\geq\tilde{\Omega}(\sqrt{\alpha d})$.
@@ 2.2 Proof via a Strong Fingerprinting Lemma
The disadvantage of using FPC as black-box for DP lower bounds is the fact that [Tardos08]’s analysis is quite involved, so it is hard to gain an end-to-end understanding of the process, and in particular, to generate hard instances in more complicated settings (e.g., exponential families [KamathMS22]). Therefore, later results simplified the construction and especially the analysis, at the cost of considering more restricted adversaries that must estimate the average of most coordinates, and not just the “marked” ones (which usually suffices for the aggregation tasks we are interested in). This led to the development of the Fingerprinting Lemma (described below) which serves as the most common tool for proving lower bounds for approximate DP algorithms.
@@
Now the same approach of [BunSU17, DworkSSUV15] yields that we can increase the correlation by taking $d=\tilde{\Theta}(n^{2})$ independent columns, [^2] An additional factor of $\log^{2}n$ is hidden inside the $\tilde{\Theta}$ due to the $\log n$ factor that we lose in our new fingerprinting lemma. However, we ignore it for the sake of this presentation as it is only a low-order term. which leads to a tracing algorithm. But unlike [BunSU17, DworkSSUV15], we obtain a much stronger “FPC-style” tracing algorithm, which enables to apply a similar PAP approach as described in Section 2.1. As a corollary that is of independent interest, the above approach results in a new fingerprinting code that has a simpler analysis than the one of [Tardos08], described in Appendix A.
@@
[…]